Juan's model racer is designed to travel at two-thirds the speed of Raul's racer. If the cars start in the same place, but travel in opposite directions, they are 75 cm apart after 5 seconds. What is the rate of each racer?
Answer by nyc_function(2741)

Let x = the speed of Raul's racer, then 2x/3 = the speed of Juan's racer
In 5 seconds the distance was 75cm. Using the distance formula.
D = (R) (T)
75 = (x + 2x/3)5
15 = (x +2x/3)
15 = (3x + 2x)/3
45 = 5x
9 = x This is Raul's racer speed
9(2/3) or 6cm = Juan's racer speed